OpenStudy (anonymous):

An instructor has graded 26 exam papers submitted by students in a class of 27 students, and the average so far is 68. (The maximum possible score is 100.) How high would the score on the last paper have to be to raise the class average by 1 point?

OpenStudy (aripotta):

68(26) + x/27 = 69 1768+x/27 = 69 1768+x = 1863 x = 95